# L-PR: Exploiting LiDAR Fiducial Marker for Unordered Low Overlap Multiview Point Cloud Registration

### v2

@@ V-B Evaluation of Point Cloud Registration Accuracy
V-C Application 1: 3D Asset Collection
An important application of point cloud registration is the 3D asset collection [34]. Thus, we evaluate the instance reconstruction quality of L-PR in this test. Data. The experimental setup is depicted in Fig. 9. Four 69.2 cm $\times$ 69.2 cm ArUco [11] markers are placed in the environment. We employ the Livox MID-40 LiDAR to scan the vehicle from five significantly different viewpoints. As shown in Fig. 9, this is also a challenging low-overlap case, with the average and minimum overlap rates being 13.39% and 1.02%, respectively. To evaluate the instance reconstruction quality, we use Supervisely [35], a popular 3D annotation tool, to extract the point cloud of the vehicle from the reconstruction result. Since the manufacture and model (Mercedes-Benz GLB) of the vehicle to be reconstructed are known, a high-fidelity 3D model acquired from a 3D assets website acts as the ground truth shape. Metrics. Following [34], we compute the Chamfer Distance and Recall between the ground truth shape and the reconstruction result. Results and Analysis. The qualitative result is shown in Fig. 9. As seen, SGHR is only able to align the third and fifth scans, while Teaser++ fails in all scenes even when pairs with overlap are manually selected and provided. By contrast, L-PR successfully aligns the unordered multiview point clouds and our reconstructed result preserves intricate details. The Chamfer Distance and Recall of our reconstructed result are 0.003 and 96.22%, respectively. This indicates [34] that L-PR can serve as a convenient tool for collecting high-fidelity 3D assets using the LiDAR sensor.
Figure 9: An illustration of the experimental setup and a comparison of reconstruction results of Ours, SGHR [1], and Teaser++ [3].
@@ V-C Application 1: 3D Asset Collection
V-D Application 2: Training Data Collection and Enhancement of Existing Learning-Based Method
Training data is crucial for learning-based methods. Most existing methods [1, 2] are trained on the 3DMatch dataset [6]. Augmenting the training data is beneficial for improving the generalization ability of learning-based methods. As demonstrated in Sections V-B and V-C, L-PR can automatically and efficiently register multiview low overlap 3D point clouds. Thus, it can serve as a valuable tool for collecting training data. In this section, we demonstrate that the training data collected using L-PR can enhance the performance of the SOTA method, SGHR [1], on various benchmarks, including 3DMatch [6], ETH [21], and ScanNet [22]. In particular, using L-PR, we process all the point clouds shown in Figs. 8 and 9, including 11 scenes with 33 scans, into the format required for training by SGHR. The newly collected data is named Livox-3DMatch. Then, we train SGHR from scratch using a mix of 3DMatch and Livox-3DMatch. In particular, the number of pairs in the training data increased from 14,400 to 17,700, an increase of 22.91%, thanks to the introduction of our data. Finally, we compare the performance of the SGHR models trained with only 3DMatch and with 3DMatch plus our data on popular benchmarks [6, 21, 22]. Metrics. Following [1], we use Registration Recall (RR) to evaluate performance on 3DMatch [6] and ETH [21], and RMSEs to report performance on ScanNet [22]. Results and Analysis. The quantitative results are presented in Table III. As seen, the RR of SGHR is increased by 2.90% on 3DMatch [6] and 4.29% on ETH [21]. The translation error and rotation error on ScanNet [22] are decreased by 22.72% and 11.19%, respectively.
V-E Application 3: Reconstructing a Degraded Scene
As previously introduced, the use of fiducial markers can enhance the robustness of reconstruction and localization in degraded scenes. Therefore, in this test, we evaluate our method in a textureless, degraded scene. Data. The setup is shown in the reference of Fig. 10. This scenario has repetitive structures and weak geometric features. We attach thirteen 16.4 cm $\times$ 16.4 cm AprilTags to the wall. The LiDAR scans the scene from 11 viewpoints. We also captured 72 images with an iPhone 13 to use as input for SfM-M [14]. The ground truth trajectories are given by an OptiTrack Motion Capture system. Competitors and Metrics. Considering the competitors [1, 3] struggle in the degraded scene, we add SfM-M [14], the state-of-the-art marker-based SfM method, as a competitor. This addition enables a comprehensive and meaningful comparison. We employ the RMSEs as the metric. Results and Analysis. The qualitative results and trajectory comparison are presented in Fig. 10. Once again, SGHR and Teaser++ struggle in this scene. The quantitative results shown in Table IV demonstrate that the proposed approach achieves better localization accuracy, which is expected given that LiDAR is a ranging sensor.
Figure 10: Visualization of the scene reconstruction and localization results of different methods.
@@ V-E Application 3: Reconstructing a Degraded Scene
SfM-M [14] 0.0551 0.0491
Ours 0.0490 0.0384
V-F Application 4: Localization in a GPS-denied Environment
Localization is also a crucial application of point cloud registration methods. Fiducial markers are a popular tool for providing localization information in GPS-denied environments, such as indoor parking lots. In this test, we compare the proposed method with KISS-ICP [4] in terms of localization accuracy. As shown in Fig. 11(a), four 69.2 cm $\times$ 69.2 cm ArUco [11] markers are deployed in the environment. The vehicle, equipped with an RS-Ruby 128-beam mechanical LiDAR, follows an 8-shaped trajectory without pausing and samples 364 LiDAR scans. We conduct the experiment on the roof of a large parking lot to acquire the ground truth trajectory from the Real-Time Kinematic. However, our method is also applicable to indoor GPS-denied environments. As seen in Fig. 11(b) and Table V, our method exhibits less drift in the middle of the trajectory.
Figure 11: (a): An illustration of the experimental setup. (b): Comparison of the trajectories given by different methods.
TABLE V: Comparison with KISS-ICP [4] regarding localization.
@@ V-F Application 4: Localization in a GPS-denied Environment
KISS-ICP [4] 0.1976 0.1617
Ours 0.1715 0.1394
V-G Application 5: 3D Map Merging
By cooperating with the algorithm introduced in [36], the proposed L-PR can be utilized for 3D map merging, as shown in Fig. 12. These 3D maps are extremely low overlap cases. In particular, the overlap rates [2] are 4.87% between map 1 and map 2, 3.96% between map 1 and map 3, and 2.59% between map 2 and map 3.
Figure 12: An illustration of 3D map merging using L-PR.
VI Conclusion
@@
In this work, we develop L-PR, a novel framework that exploits LiDAR fiducial markers for unordered low overlap multiview point cloud registration. The markers in this work are thin sheets of paper attached to other surfaces, without impacting the 3D environment. L-PR consists of an adaptive threshold marker detection method and two levels of graphs. The adaptive threshold method robustly detects markers even when there are dramatic changes in viewpoints among scans, thus overcoming the limitation of [19]. The proposed two-level graph efficiently aligns the low overlap multiview point clouds directly from an unordered set by optimizing a MAP problem. We conduct qualitative and quantitative experiments to demonstrate the superiority of L-PR over previous SOTA methods and the diverse applications of L-PR, including 3D asset collection, training data collection, reconstruction of a degraded scene, localization in a GPS-denied environment, and 3D map merging. In particular, we collect a new training dataset named Livox-3DMatch using L-PR. The Livox-3DMatch dataset expands the original 3DMatch training data from 14,400 pairs to 17,700 pairs, representing a 22.91% increase. Training on this augmented dataset improves the performance of the SOTA learning-based method SGHR [1] by 2.90% on 3DMatch [6], 4.29% on ETH [21], and 22.72% (translation) / 11.19% (rotation) on ScanNet [22]. Future work includes utilizing L-PR to collect more 3D assets and point cloud registration training data to benefit the development of learning-based methods [34, 1].

### v3

@@ VI-C Application 1: 3D Asset Collection from Sparse Scans
MDGD [9] 94.26 96.06 0.49/0.44 20.51/19.82
MDGD [9] + Livox-3DMatch (our data) 95.97 98.95 0.38/0.31 18.91/17.36
Results and Analysis. The qualitative results are shown in Fig. 9 and Fig. 10. As shown in Fig. 9, MDGD [9] and SGHR [1] successfully align the third and fifth scans. In particular, the third and fifth scans have the highest overlap ratio of 59.33% among all pairs as they are captured from similar perspectives. However, when dealing with point clouds that have lower overlap ratios, the competitors [9, 1, 7, 8, 3] struggle to register them. This comparison illustrates that, although these existing methods can handle some unseen scenarios with high overlap, they are not generalizable to novel low overlap cases. Moreover, the failure of these existing methods in the instance reconstruction task implies that they are not suitable for efficient 3D asset collection. By contrast, the proposed method successfully registers the unordered multiview point clouds. Especially, as presented in Fig. 10, our reconstructed shape preserves intricate details well compared to the ground truth shape. From a quantitative perspective, the Chamfer Distance and Recall of our reconstructed result are 0.003 and 96.22%, respectively. The decent performance in terms of reconstruction quality metrics [25] indicates that L-PR can serve as a convenient, efficient, and low-cost tool for collecting high-fidelity 3D assets using the LiDAR sensor. In particular, we consider the proposed method efficient, as it only requires four or five scans from dramatically changed viewpoints to reconstruct the complete shape.
@@
VI-D Application 2: Training Data Collection and Enhancement of Existing Learning-Based Methods
Training data is crucial for learning-based methods. Most existing methods [1, 9] are trained on the 3DMatch dataset [6]. Augmenting the training data is beneficial for improving the generalization ability of learning-based methods. However, the existing methods cannot be utilized for collecting training data in unseen scenes, as unseen scenes imply out-of-distribution cases and the limited effectiveness of the existing methods. The proposed method can serve as a valuable tool for collecting training data. In this test, we demonstrate that training data collected using L-PR can enhance the performance of the SOTA methods across various benchmarks.
Data. Using the proposed method, all the point clouds shown in Figs. 8 and 9, comprising 11 scenes with 33 scans, are aligned and processed into the format required for training by MDGD [9] and SGHR [1]. The newly collected data is named Livox-3DMatch. We train the models [9, 1] from scratch using only 3DMatch and a combination of 3DMatch and Livox-3DMatch.
Benchmarks and Metrics. We compare the performance of the models [9, 1] trained with and without our data on three popular benchmarks: 3DMatch [6], ETH [23], and ScanNet [24]. Following [1], we use Registration Recall (RR) to evaluate performance on 3DMatch [6] and ETH [23], and RMSEs to report performance on ScanNet [24].
Results and Analysis. First, the number of pairs in the training data increases from 14,400 to 17,700, a boost of 22.91%, thanks to the introduction of our data. The quantitative results are reported in Table III. As seen, the RR of SGHR [1] is increased by 2.90% on 3DMatch [6] and 4.29% on ETH [23]. The translation error and rotation error on ScanNet [24] are decreased by 22.72% and 11.19%, respectively. The RR of MDGD [9] is improved by 1.71% on 3DMatch [6] and 2.89% on ETH [23]. The translation error and rotation error on ScanNet [24] are decreased by 22.45% and 7.80%, respectively. The reason behind these improvements is mentioned in Section V: the introduction of our data enriches the learnable features during training, which benefits the generalizability of the models.
Figure 11: An illustration of the experimental setup and a visual comparison of the proposed method against the SOTA methods (MDGD [9], SGHR [1], SE3ET [7], GeoTrans [8], Teaser++ [3], and SfM-M [16]) in a degraded scene.
Figure 12: A comparison of the sensor trajectories obtained from the proposed method and SfM-M [16]. G.T. refers to the ground truth.
VI-E Application 3: Reconstructing a Degraded Scene
An important application of fiducial markers in the real world is enhancing the robustness of reconstruction and localization in degraded scenes when additional sensors are unavailable. Therefore, in this test, we evaluate our method in a textureless, degraded scene.
Data. The setup is shown in the reference of Fig. 11. This scenario has repetitive structures and weak geometric features. We attach thirteen 16.4 cm $\times$ 16.4 cm AprilTags to the wall. The LiDAR scans the scene from 11 viewpoints. We also captured 72 images with an iPhone 13 to use as input for SfM-M [16]. The ground truth trajectories are given by an OptiTrack Motion Capture system.
Competitors and Metrics. The competitors are the SOTA point cloud registration methods [9, 1, 7, 8, 3]. However, considering these competitors struggle in the degraded scene, we add SfM-M [16], the SOTA marker-based SfM method, as a competitor. This addition enables a comprehensive and meaningful comparison. We employ the RMSEs as the metric.
Results and Analysis. The qualitative results are presented in Fig. 11. As seen, none of the point cloud registration methods [9, 1, 7, 8, 3] can align a single pair of point clouds. This is because these cases only have planar overlap regions and thus are even more challenging than those in Fig. 8 and Fig. 9. In comparison, the proposed method and SfM-M [16] successfully reconstruct this degraded scene as they utilize the fiducial markers. The comparison of these two methods in terms of sensor trajectories is shown in Fig. 12. The quantitative results shown in Table IV demonstrate that the proposed approach achieves better localization accuracy, which is expected given that LiDAR is a ranging sensor.
@@ VI-E Application 3: Reconstructing a Degraded Scene
SfM-M [16] 0.055 0.049
Ours 0.049 0.038
VI-F Application 4: Localization in a GPS-denied Environment
Localization is also a crucial application of point cloud registration methods. Fiducial markers are a popular tool for providing localization information in GPS-denied environments, such as indoor parking lots. In this test, we evaluate the proposed method in this context.
Data. The experimental setup is shown in Fig. 13(a): four 69.2 cm $\times$ 69.2 cm ArUco [13] markers are deployed in the environment. The vehicle, equipped with an RS-Ruby 128-beam mechanical LiDAR, follows an 8-shaped trajectory without pausing and samples 364 LiDAR scans. We conduct the experiment on the roof of a large parking lot to acquire the ground truth trajectory from the Real-Time Kinematic.
Competitors and Metrics. We compare the proposed method with KISS-ICP [4] in terms of localization accuracy. We employ the RMSEs as the metric.
Results and Analysis. The visual comparison and quantitative comparison of localization results are presented in Fig. 13(b) and Table V, respectively. Our method exhibits less drift in the middle of the trajectory and demonstrates better overall localization accuracy.
Figure 13: (a): An illustration of the experimental setup. (b): A comparison of the trajectories given by the proposed method and KISS-ICP [4]. G.T. refers to the ground truth.
TABLE V: Comparison of localization results between KISS-ICP [4] and our method.
@@ VI-F Application 4: Localization in a GPS-denied Environment
KISS-ICP [4] 0.198 0.162
Ours 0.172 0.139
VI-G Application 5: 3D Map Merging
To validate the performance of the proposed method in large-scale scenarios, we apply L-PR to the 3D map merging task in this test, which involves merging multiple large-scale, low overlap 3D maps into a single frame.
Data. We collected three large-scale LiDAR maps. They are constructed using the SOTA LiDAR-based SLAM method, Traj-LO [47], by scanning the York University campus with a Livox MID-40 LiDAR. The ground truth poses between the maps are manually obtained using CloudCompare [44].
Competitors and Metrics. The competitors are SOTA multiview point cloud registration methods, including MDGD [9] and SGHR [1]. Moreover, unlike previous tests, the LiDAR fiducial markers on the 3D maps are localized using the algorithm from our previous work [48], and the marker detection results serve as input for the L-PR pipeline. We employ the RMSEs as the metric.
Results and Analysis. The visual and quantitative comparisons are shown in Fig. 14 and Table VI, respectively. As shown in Fig. 14, neither MDGD nor SGHR can address this challenging task. This is due to the fact that the overlap in large-scale maps is too scarce. Specifically, the overlap rates [2] are 4.87% between map 1 and map 2, 3.96% between map 1 and map 3, and 2.59% between map 2 and map 3. On the other hand, both MDGD and SGHR start by analyzing the features of each individual point cloud. As the scale of the point cloud becomes larger, the portion of the features belonging to the overlap regions decreases, making them unsuitable for large-scale, low overlap map merging. In addition, as the scale of point clouds becomes larger, the absolute error values of MDGD and SGHR also increase. In comparison, even though the scales of the point clouds in this test are much larger than those in the previous test, the proposed method successfully merges these large-scale, low overlap 3D maps. This stems from the observation that the proposed method focuses on utilizing the thin-sheet LiDAR fiducial markers and is not sensitive to scale changes. As introduced in [48], the accuracy of marker localization degrades as the scale increases. Thus, the absolute error values of our method also increase compared to previous tests.
Figure 14: A comparison of the 3D map merging results of MDGD [9], SGHR [1], and our method.
TABLE VI: Comparison of 3D map merging results between MDGD [9], SGHR [1], and our method.
[…]
MDGD [9] 2.7883 1.9210
SGHR [1] 4.0933 2.3925
Ours 0.2305 0.1771
VI-H Ablation Studies
The proposed L-PR framework is composed of two levels of graphs. To demonstrate the necessity of this overall architecture, we conduct ablation studies in this section. In particular, we study the effects of removing the first and second graphs in two cases: Fig. 9 and the kitchen scene in Fig. 8. The visual comparison and quantitative results are shown in Fig. 15 and Table VII, respectively. When the first graph is removed, the factors representing the relative poses between frames in Fig. 5 have to be set to identity due to the lack of initial values provided by the first graph. However, the role of the second graph is to further optimize the variables based on their initial values rather than finding the optimal solution from scratch. Consequently, as shown in Fig. 15 and Table VII, severe misalignment or degradation in registration accuracy occurs. When the second graph is removed, the multiview point clouds are directly registered using the initial values obtained from the first graph, without any further refinement. As shown in Fig. 15 and Table VII, removing the second graph causes a slight degradation in registration accuracy. Moreover, the degradation in the kitchen case of Fig. 8 is slighter than that in Fig. 9. This is because the effect of the second graph is case-by-case, determined by the quality of the initial values provided by the first graph. Namely, the better the initial values are, the less important the second graph becomes. However, in the real world, pose estimation of markers cannot be perfect. In addition, as seen in Fig. 5, we also add the marker corners to the second graph so that the pose estimation of each individual marker can be further optimized along with other variables in the graph. Therefore, it is beneficial to apply the second graph in practice. In summary, L-PR adopts a coarse-to-fine pipeline, where the first graph corresponds to the coarse stage, while the second graph is the fine stage. Removing any of them will cause performance degradation.
Figure 15: A comparison of the point cloud registration results without the first graph, without the second graph, and with both graphs for two cases.
TABLE VII: Ablation studies on the first and second graphs.
Scene Framework $\mathrm{RMSE}_{T}$ (m) $\downarrow$ $\mathrm{RMSE}_{R}$ (rad) $\downarrow$
Fig. 9 without first graph 0.431 0.459
without second graph 0.078 0.084
with both graphs 0.066 0.072
Fig. 8 without first graph 0.088 0.112
without second graph 0.021 0.073
(Kitchen) with both graphs 0.019 0.067
VI-I Limitations
Despite the promising results of L-PR, there are potential limitations. Firstly, although LiDARs are robust to unideal illumination conditions, adverse weather phenomena such as rain, snow, and fog can affect LiDAR measurements, thereby reducing the effectiveness of L-PR. Secondly, the adopted low-cost LFMs, made of thin-sheet paper or boards, may deform after long-term use in the wild. However, this is not a concern for one-time applications such as data collection. Finally, deploying the LFMs requires some labor, but their value is demonstrated in this paper.
VII Conclusion
@@
In this work, we develop L-PR, a novel framework that exploits LiDAR fiducial markers for unordered low overlap multiview point cloud registration. The markers in this work are thin sheets of paper attached to other surfaces, without impacting the 3D environment. L-PR consists of an adaptive threshold marker detection method and two levels of graphs. The adaptive threshold method robustly detects markers even when there are dramatic changes in viewpoints among scans, thus overcoming the limitation of [21]. The proposed two-level graph efficiently aligns the low overlap multiview point clouds directly from an unordered set by optimizing a MAP problem. We conduct qualitative and quantitative experiments to demonstrate the superiority of L-PR over the state-of-the-art competitors and the diverse applications of L-PR, including 3D asset collection, training data collection, reconstruction of a degraded scene, localization in a GPS-denied environment, and 3D map merging. In particular, we collect a new training dataset named Livox-3DMatch using L-PR. The Livox-3DMatch dataset expands the original 3DMatch training data from 14,400 pairs to 17,700 pairs, representing a 22.91% increase. Training on this augmented dataset improves the performance of the SOTA learning-based methods on various benchmarks. In particular, the RR of SGHR [1] is increased by 2.90% on 3DMatch [6] and 4.29% on ETH [23]. The translation error and rotation error on ScanNet [24] are decreased by 22.72% and 11.19%, respectively. The RR of MDGD [9] increases by 1.71% on 3DMatch [6] and 2.89% on ETH [23]. The translation error and rotation error on ScanNet [24] decrease by 22.45% and 7.80%, respectively. Future work includes utilizing L-PR to collect more 3D assets and point cloud registration training data to benefit the development of learning-based methods [9, 1, 49].